OpenStudy (anonymous):

Jeanette wants to tile the floor of a room in her house. The square tiles measure 3/4 ft on each side. The room is 10 ft wide. a. Write an inequality to describe how many tiles are needed to make one row of tiles across the width of the room. b. Solve the inequality. c. How many tiles should Jeanette buy to form one row?

OpenStudy (anonymous):

x= the number of tiles.\[10\le 3/4*x<11\] x is a positive integer. 3/4*x=10 x=10(4/3) x=40/3 x=13 1/3 tiles Since x must be a positive integer (they don't sell partial tiles at the store) and must cover a distance of 10 feet, then she will have to buy a total of 14 tiles.